Alan bought $20.40 worth of cheese and butter. A pound of cheese cost $2.40 and a pound of butter cost $3.20 The cheese and butter together weighed 7.5 pounds. How many pounds of each did he buy?
Answer by CubeyThePenguin(3113)

c = pounds of cheese
b = pounds of butter

c + b = 7.5
2.4c + 3.2b = 20.40

Multiply the first equation by 3.2.

3.2c + 3.2b = 24
2.4c + 3.2b = 20.40

So, 0.8c = 3.6 --> c = 4.5.

Alan bought 4.5 pounds of cheese and 3 pounds of butter.